This table indicates how rapidly water evaporates from an open 720-gallon tank.

Days/ Remaining Water(gallons)
0/ 720
10/ 660
20/ 600
30/ 540

How much water is left at the end of 90 days?

*I know that over 10 days 60 gallons of water evaporates, but I just can't find out the equation. I worked it out and got 180 gallons by just subtracting 60 from 540 seven times. I'm just confused on the equation to use for the problem.*

Answer by checkley71(8403)

THE FORMULA IS:
720-9*60=720-540=180 AMOUNT LEFT AFTER 90 DAYS.